Find the payment that should be used for the annuity due whose future value is given. Assume that the compounding period is the same as the payment period. $21,000​; quarterly payments for 10 years; interest rate 8.1​%
Answer #1

Periodic Quarterly Payment = $21,000

Interest rate = 10% compounding quarterly

Calculating the Future Value :-

Future Value= C*\frac{[(1+r)^{n}-1]}{r}

Where, C= Periodic Payments = $21000

r = Periodic Interest rate = 8.1%/4 = 2.025%

n= no of periods =10yrs*4 = 40

Future Value= 21,000*\frac{[(1+0.02025)^{40}-1]}{0.02025}

Future Value= 21,000*\frac{[2.2297908953-1]}{0.02025}

Future Value = $1,275,338.71